# See pages - of preprint.docx.pdf

### v2

@@
See pages - of preprint.docx.pdf
@@
Reconfigurable legged metamachines that run on autonomous modular legs
Chen Yu, David Matthews Jingxian Wang, Jing Gu, Douglas Blackiston, Michael Rubenstein, Sam Kriegman
Co-first author.
@@
Supplemental Movies S1 and S2 can also be streamed online and are linked on our project page:
@@
Metamachines in the wild. Reconfigurable legged metamachines behave autonomously across sand, mud, grass, tree roots, plant litter, mulch, gravel, bricks, concrete, and combinations thereof.
@@
Acrobatic behaviors and resilience to damage. Reconfigurable metamachines behave, resist adversarial perturbations, and adapt to damage. Control policies use internal sensing only; motion capture was not supplied to the policy, it is used only for the behavioral analyses in Figs. 2, 3 and 5 in the manuscript, as well as that of Figs. S1 and S2 below.
@@
The two links and central sphere of each module were 3D printed with PAHT-CF filament on a Bambu Lab X1C 3D Printer. The PCB cradle was printed with PLA filament. Table S1 lists 3D printing parameters.
@@
Each module has sufficient onboard processing capabilities for fully autonomous closed-loop control. However, for convenience we employ a remote computer to calculate module actions. Each module is connected to a remote computer through WiFi. The module publishes all the sensor data and a unique module identification number (bound to MAC address) to the computer over UDP at a frequency of 100 Hz. Sensor data used for observations include: IMU data and motor encoder data. Module reports additional sensing information (e.g. voltage, current, motor temperature, etc.) for logging purposes. The remote computer publishes motor position targets and PD control parameters $K_{p}$ and $K_{d}$ to each module via UDP at 20Hz. Since the modules publish sensor data faster than the control frequency, the remote computer only utilizes the most recent data.
@@
We add a cosine mapping to the joint position to deal with the general case where the joint can rotate continuously. All other details of the observation space are described as part of the Methods in the manuscript (under the heading, “Observation Space”).
@@
For rolling, the reward function is defined by two terms: forward reward and action rate penalty. To compute the forward reward, we define $\psi$ to be the angle between a module’s two links, and $\bm{\ell}$ is defined as the line which bisects $\psi$. The line perpendicular to $\bm{\ell}$ which intersects the center of the sphere and is parallel to the plane formed by the two links is denoted $\bm{\ell}^{\prime}$. We define the angular velocity ${\omega}_{f}$ that the module rolls forward as
@@
the dot product of the angular velocity $\omega$ in the body’s frame and the vector $\bm{\ell}^{\prime}$. Given a desired forward rolling speed, $\omega^{\ast}$, the rolling reward, $r_{\text{roll}}$, is calculated as:
@@
We penalize the action rate of all modules in a metamachine in simulation to encourage the smoothness of actions and help the sim-to-real transfer, as this term serves as a regulator in the policy optimization. The action rate reward at time step $t$ is defined as:
@@
[…]
@@
[…]
@@
We use CrossQ  (50) for training both single modules and multi-module metamachines. A relatively small policy network (256, 256) was used for sample efficiency alongside a larger critic network (1024, 1024). Policies were trained $10^{6}$ time steps.
@@
Since we desire metamachines composed of these modules to perform powerful, athletic behaviors, inter-module connections through docks must be strong enough to endure stress in all directions. We thus designed the docks to be honeycomb-shaped and non-flat so that the external force and torque can be largely combated by the normal force on the docking surface and the combined torque of the normal forces. This design helps to relieve stress on connecting bolts and bolts’ contact points on the metamachine, reducing failures and increasing rigidity of the docks.
@@
With the design choices described above, our modules can dock together to form a wide variety of metamachines. Consider two modules, $i$ and $j$, which are connected together via a dock to form a metamachine. We first assume that these two modules are not identical. Because any two docks can connect together, each module has 18 docking points, and docks have 3-fold rotational symmetry, we have $18\times 18\times 3$ possible configurations. However, due to interference, two parallel modules cannot dock through docks on the sides of links while remaining parallel. Each module has 12 such docking points, and so interference eliminates $12\times 12$ configurations.
@@
In reality, the modules are identical, so a metamachine formed by module $i$ and $j$ will be the same if module $i$ takes module $j$’s place and vice versa. As such, we need to divide the numbers we have obtained by 2 to count the number of unique configurations. However, if both modules are connected through the same dock, swapping $i$ and $j$ will result in an identical configuration even if the modules are not identical, and these cases should not be divide by 2. Each module has 18 docking points, and docks have 3-fold rotational symmetry, thus we have $18\times 3$ configurations where both modules are connected through the same dock. However, among these possibilities, 12 cases will lead to interference as explained above and should be left out. So, in total we have $18\times 3-12$ configurations where both modules are connected through the same dock. After accounting for these cases, in total there are $(18\times 18\times 3-12\times 12+18\times 3-12)/2=435$ unique two module metamachines.
@@
As the number of modules increases, interference between modules and the symmetry of the metamachine becomes much more challenging to evaluate, so we will provide an estimation of the number of unique configurations instead. Let’s assume that there are $M$ unique configurations for an $N$ module metamachine. When adding another module to the $N$ module metamachine, we have $216N+$ 2 free docks on an $N$ module metamachine and 18 docks on the added module, giving us a total of $(16N+2)\times 18\times 3\times M\approx 864NM$ possible configurations. Considering that swapping the added module with any module in the $N$ module metamachine will result in an identical metamachine, we approximately have $864NM/(N+1)$ unique metamachines. Using mathematical induction, we get a good estimate for the number of unique $N$ module metamachines to be $864^{(N-1)}/N$. This number would grow to approximately $10^{11}$ when $N=5$.
@@
To find good metamachine configurations with high locomotive ability, we encoded the combinatorial configuration space into a continuous 8D latent space using a variational autoencoder (VAE). A VAE was used because it transforms the search problem into a continuous one, which simplifies optimization (52), and because we have considerable experience using this particular encoding. We searched for good configurations within this compressed latent design space using Bayesian Optimization (BO). We used multilayer perceptrons (MLP) for both the VAE encoder and decoder networks. The size of each hidden layer of the VAE encoder was (512, 128, 64, 64, 16) and the size of each hidden layer of the VAE decoder was (64, 64, 128, 512).
S4.3 Representing different metamachine configurations (VAE input)
To input the topologies of metamachines into a VAE, we represent an assembled metamachine as a tree, where each node is an attached module. Each configuration tree can be uniquely represented as a sequence of integers with each connection pair represented by four items: the module ID of the parent module, the position ID of the dock on the parent module, the dock ID of the child module, and the orientation ID of these two docks. In this way, an assembled metamachine of $N$ modules can be represented by $4\times(N-1)$ integers. To represent metamachines with $M$ modules when $M<N$, the final $4\times(N-M)$ integers are set to a reserved value which denotes “not present”. We employed a one-hot encoding on each integer in the sequence, and then fed the concatenated result to the VAE.
@@
Because training time can vary greatly when evaluating configurations with differing numbers of modules, we used asynchronous Bayesian Optimization (BO), allowing multiple workers to comb the latent space for good configurations, in parallel asynchronously. We use the method in  (31) to penalize the acquisition function using information about configurations that are still under evaluation. BO was used to identify configurations that were able to locomote well. The fitness of a metamachine for the purpose of BO was taken to be the average accumulated reward across the final 10% of episodes. Candidate designs identified through this method underwent training on additional behaviors before being selected for assembly.
@@
S5 Metamachine controller
@@
Inspired by  (53), the reward of the walking task is a linear combination of six terms: forward velocity tracking reward $r_{\text{forward}}$, angular velocity tracking reward $r_{\text{walk-turn}}$, action rate penalty $r_{\text{action}}$, joint falling penalty $r_{\text{fall}}$, motor velocity penalty $r_{\text{motor-vel}}$, and joint acceleration penalty $r_{\text{motor-acc}}$. The forward velocity tracking reward is calculated by
@@
[…]
@@
where ${\omega}^{\ast}_{z}$ is the desired angular velocity around the global z axis, which is set to 0 here, and the ${\omega}_{z}=\bm{\omega}\cdot\bm{g}_{p}$ is the z component of angular velocity in the global frame; $\bm{\omega}$ is the metamachine’s angular velocity and $\bm{g}_{p}$ is the projected gravity. The hyperparameter $\sigma_{\text{walk-turn}}$ is set to 0.15. The action rate penalty is calculated in the same way as training for the single module as in Eqn. S3. The joint falling penalty $r_{\text{fall}}$ is the number of joints (the sphere part of the module) touching the floor for each time step. The joint velocity penalty is calculated by
@@
which will be clipped from 00 to $1\times 10^{5}$; $\dot{{\theta}}_{i}$ is the joint velocity of joint $i$ and $\dot{\theta}_{\text{limit}}$ is the joint velocity limit whose value is 10 radians/s in our experiment. The joint acceleration penalty is calculated by
@@
Similar to the single module training, the action output by the policy, $\bm{a}_{t}$, is offset by predefined values for each joint. Before feeding the offset action to the PD controller, we filter the action with a Butterworth filter (high cutoff frequency is 3; low cutoff frequency is 0; filter order is 2) to promote the smoothness of the actions  (54). The action is then clipped to ±$\pm 1.2$ radians and sent to the PD controller. The torque output by the PD controller is again clipped by the TN curve of the real motor.
@@
Although we have a fixed maximum episode length, we terminate each episode earlier if and when the metamachine falls over. Inspired by  (54), falling is determined by the deviation of the body’s upward position $d$, which is defined by:
@@
$r_{\text{pose}}=r_{\text{upward}}\,r_{\text{joint}}\$ (S12)
where $r_{\text{upward}}$ measures how well the metamachine stays in an upright position, defined as $r_{\text{upward}}=d$, where $d$ is the deviation of the body’s upward position defined in Eqn. S11. The pose reward $r_{\text{joint}}$ measures how close the current joint positions $\bm{\theta}$ are to the initial joint positions $\bm{\theta}_{0}$:
@@
The action rate penalty is calculated as in Eqn. S3. The weight for the pose reward term is 1 and the weight for the action rate is $-0.02$. The hyperparameter $\sigma_{\text{joint}}$ here is set to 10.
@@
To achieve dynamic self-righting behavior, we increase the action clipping from ±$\pm 1.2$ to ±$\pm 3.14$ radians.
@@
The observation space is the observation space for walking plus an additional dimension for a jumping command dimension, similar to  (55). The jumping command is set to 0 except when the metamachine is commanded to jump. In simulation, the jumping command is randomly sampled as 0 or 1 every 100 timesteps during training and is automatically reset to 0 when the metamachine reaches the desired height. After deploying the policy in the real world, the jump command is nominally 0 and can be set to 1 by a user. Since we assume only proprioceptive sensing on the physical hardware, rather than using a state estimator as in  (55) to measure the height of the metamachine for resetting the command, we set a fixed activation period $T_{\text{activate}}$ for the jumping command. $T_{\text{activate}}$ is set to 0.75 seconds in our experiment.
@@
The rewards of the jumping task consist of five terms: upward reward $r_{\text{upward}}$, pose reward $r_{\text{pose}}$, height track reward $r_{\text{height}}$, jump bonus $r_{\text{jump}}$, and turn reward $r_{\text{jump-turn}}$. The upward reward $r_{\text{upward}}$ is to prevent the metamachine from falling over. It is defined in the same way as in Eqn. S12. The pose reward $r_{\text{pose}}$ is to encourage the metamachine to retain the joint angles from the initial pose when the jumping command is not triggered. The pose reward is set to 0 when the jump signal is triggered (the jumping command in the observation is set to 1); when the jump signal is not triggered, the pose reward is defined as in Eqn. S12. The height track reward is to encourage the metamachine to jump to a desired height when the jumping command is triggered. When the jump signal is not triggered, $r_{\text{height}}=0$, otherwise
@@
where $h$ is the height of the torso module and $h^{\ast}$ is the desired jumping height. The values of $h^{\ast}$ for each metamachine can be found in Table S4. To further encourage the metamachine to jump, a bonus $r_{\text{jump}}$ is used, which is set to 1 when the jump signal is triggered and the metamachine is not touching the ground; otherwise $r_{\text{jump}}=0$. The turn reward $r_{\text{jump-turn}}$ is to encourage the metamachine to turn when jumping. When the jump signal is triggered,
@@
[…]
@@
To achieve dynamic jumping actions, we increase the action clipping to ±$\pm 2.5$ radians from ±$\pm 1.2$ radians for walking.
@@
We model the amputation-agnostic control problem as a POMDP. We desire a single policy that can control a metamachine after arbitrary amputations. We define the state space of the metamachine $\mathcal{S}_{\text{meta}}$ as the collection of the state space of each module $\mathcal{S}_{\text{module}}$ in this system:
@@
where $\mathcal{S}_{\text{module}}^{i}$ is the state space of the $i$-th module and $N$ is the maximum number of modules in the system. The state space of each module includes projected gravity, angular velocity, cosine of the joint position, and the joint velocity. The action of this POMDP is the joint position of each module, and the reward function varies across the different reduced configurations (example amputations) used in training the amputation-agnostic policy (a more detailed explanation is provided below). We further reduce this POMDP to a sequence modeling problem  (51). Similar to  (56), we use a Transformer model  (57) to distill a general controller from several amputation-specific teacher controllers. However, rather than using a Transformer encoder as a student policy to directly map states to actions, we use a causally masked transformer decoder to fit a sequence of states and actions. By conditioning on the current state $\mathcal{S}_{\text{meta}}$, the transformer should predict the optimal action.
@@
For a time step $t$, we feed the most recent $K$ timesteps ($K=60$ in the experiment) of module states and actions into the Transformer, for a total of $K\times(N+1)$ tokens. Each of the module states and actions will be fed into its dedicated linear layer with layer normalization. We then add a positional embedding to the embedding of each token. Similar to Decision Transformer  (51), rather than using a standard positional embedding, we use the embedding of each timestep as the positional embedding (each of the $N+1$ tokens thus have the same positional embedding). The embeddings are then fed to a Generative Pre-trained Transformer (GPT) model  (58), which predicts the next action tokens through autoregressive modeling. The hyperparameters of the GPT model are shown in Table S9.
@@
A dataset of offline trajectories is given to train this model. In our experiment, we use the training dataset in Table S7. To generate this dataset, scenario-specific expert policies were trained from scratch to control the remnant structures. For training remnant configurations with more than one module remaining, we use the same observation space, action space, and reward function as the walking policy described above. For training a single module, we use the same observation, reward, and action as the single module rolling policy described above. However, when training against amputation examples, we add more randomization to the environment compared to the standard training methods described above. In the undamaged scenario, we randomize the length of all the limbs: Each link touching the floor is either with its nominal length or a half of its nominal length. In the four-remaining-modules scenario (one limb is removed), the farthest link of the limb diagonally across from the amputated limb is randomly assigned either its nominal length or 1.5 times its nominal length. For example, if the configuration is that the left forelimb is amputated, the right hind limb will have two possible lengths: the nominal length or an extended length. This encourages the metamachine to recover to an upright position. Training settings for single limb amputations of the quadruped are shown in Table S8. In the three-remaining-modules scenario (two limbs removed), we randomize the length of the forelimbs and the torso: Each link touching the ground is either with its nominal length or a half of its nominal length. When one module remains post-amputation, the left or right link is half of its nominal length. These remnant single modules were also trained on uneven terrain featuring obstacles of random heights. The heights of the obstacles are sampled from 0 to 0.03 m.
Note that the observation space of the POMDPs for generating the training dataset is different from the state space $\mathcal{S}_{\text{meta}}$. When collecting the rollouts after these policies are trained, the state of the metamachine $\bm{s}_{\text{meta}}\in\mathcal{S}_{\text{meta}}$ is collected rather than the observation of the specific amputation example. In addition, although the different amputation-specific policies have different zero positions for joints, we normalize them to have the same zero position when recording the rollout trajectories after training them. The training of each metamachine takes $1\times 10^{6}$ timesteps. To avoid the metamachine getting stuck at a local optimum and stopping motion, we set the weight of the forward term in the reward function as 1 at the beginning of training and tuned it down to 0.8 at the time step of $2\times 10^{5}$.

### v3

# arXiv:2505.00784v3[cs.RO] 27 Feb 2026

@@
arXiv:2505.00784v3[cs.RO] 27 Feb 2026
See pages - of Manuscript.pdf
@@
Agile legged locomotion in reconfigurable modular robots
Chen Yu1, David Matthews2, Jingxian Wang2, Jing Gu, Douglas Blackiston, Michael Rubenstein, Sam Kriegman
1,2Co-first authors.
2Order determined by coin toss.
@@
Supplemental Movies S1–S3 can also be streamed online and are linked on our project page:
@@
Reconfigurable robots in the wild. Reconfigurable legged robots behave autonomously across sand, mud, grass, tree roots, plant litter, mulch, gravel, bricks, concrete, and combinations thereof.
@@
Acrobatic behaviors and resilience to damage. Robots behave, resist adversarial perturbations, and withstand body-altering damage. Control policies use internal sensing only; motion capture was not supplied to the policy, it is used only for the behavioral analyses in Figs. 2, 3 and 5 in the manuscript, as well as that of Figs. S1 and S2 below.
@@
The two links and central sphere of each module were 3D printed with PAHT-CF (Polyamide High-Temperature Carbon Fiber) filament on a Bambu Lab X1C 3D Printer. The PCB (Printed Circuit Board) cradle was printed with PLA (Polylactic acid) filament. Table S1 lists 3D printing parameters.
@@
Each module has sufficient onboard processing capabilities for fully autonomous closed-loop control. However, for convenience we employ a remote computer to calculate module actions. Each module is connected to a remote computer through WiFi. The module publishes all the sensor data and a unique module identification number (bound to MAC address) to the computer over UDP at a frequency of 100 Hz. Sensor data used for observations include: IMU data and motor encoder data. Module reports additional sensing information (e.g. voltage, current, motor temperature, etc.) for logging purposes. The remote computer publishes motor position targets and PD (Proportional-Derivative) control parameters $K_{p}$ and $K_{d}$ to each module via UDP at 20Hz. Since the modules publish sensor data faster than the control frequency, the remote computer only utilizes the most recent data.
@@
We add a cosine mapping to the joint position to deal with the general case where the joint can rotate continuously. All other details of the observation space are described as part of the Materials and Methods in the manuscript (under the heading, “Observation Space”).
@@
[…]
@@
[…]
@@
We penalize the action rate of all modules in a robot in simulation to encourage the smoothness of actions and help the sim-to-real transfer, as this term serves as a regulator in the policy optimization. The action rate reward at time step $t$ is defined as:
@@
is the fraction of the desired turning speed, capped at 1. $\omega_{\text{turn}}^{\ast}=3\,\text{rad/s}$. The final reward are calculated as:
@@
Control policies were trained simulation using the Mujoco simulator  (?). Domain randomization was employed to ensure behaviors optimized in simulation transferred with sufficient fidelity to the physical system. Mass distribution, mass, module geometry, friction, and motor parameters were randomized during training. Mass distribution was randomized by sampling the motor position offset and motor mass. Domain randomization parameters are shown in Table S2, where $m$ is the mass of the agent (single module or multi-module robot). We uniformly sample these values at each episode of the training process. We also add Gaussian noise sampled from $\mathcal{N}(0,\,0.2)\,$ to the observations and noise from $\mathcal{N}(0,\,0.1)\,$ to the actions. One of the sim-to-real transfer challenges was the latency of the WiFi communication between the remote computer that runs the RL policy and the module. To model this latency in simulation, the simulated module randomly switches between executing the current action output by the policy or the previous action, with equal probability.
@@
We use CrossQ  (?) for training both single modules and multi-module robots. CrossQ adopts the same overall update rules as Soft Actor-Critic  (?), with separate updates for the critic, the actor, and the entropy temperature. The critic is updated by minimizing the temporal-difference error between the predicted Q-value and a bootstrap target; the actor is updated to favor actions that achieve high Q-values; and the temperature parameter is adjusted to keep the policy entropy close to a target value. See Bhatt et al. (2024;  (?)) for more details. A relatively small policy network (256, 256) was used for sample efficiency alongside a larger critic network (1024, 1024). Policies were trained $10^{6}$ time steps.
@@
Since we desire robots composed of these modules to perform powerful, athletic behaviors, inter-module connections through docks must be strong enough to endure stress in all directions. We thus designed the docks to be honeycomb-shaped and non-flat so that the external force and torque can be largely combated by the normal force on the docking surface and the combined torque of the normal forces. This design helps to relieve stress on connecting bolts and bolts’ contact points on the robot, reducing failures and increasing rigidity of the docks.
@@
With the design choices described above, our modules can dock together to form a wide variety of morphologies. Consider two modules, $i$ and $j$, which are connected together via a dock to form a two-module design. We first assume that these two modules are not identical. Because any two docks can connect together, each module has 18 docking points, and docks have 3-fold rotational symmetry, we have $18\times 18\times 3$ possible configurations. However, due to interference, two parallel modules cannot dock through docks on the sides of links while remaining parallel. Each module has 12 such docking points, and so interference eliminates $12\times 12$ configurations.
@@
In reality, the modules are identical, so a design formed by module $i$ and $j$ will be the same if module $i$ takes module $j$’s place and vice versa. As such, we need to divide the numbers we have obtained by 2 to count the number of unique configurations. However, if both modules are connected through the same dock, swapping $i$ and $j$ will result in an identical configuration even if the modules are not identical, and these cases should not be divide by 2. Each module has 18 docking points, and docks have 3-fold rotational symmetry, thus we have $18\times 3$ configurations where both modules are connected through the same dock. However, among these possibilities, 12 cases will lead to interference as explained above and should be left out. So, in total we have $18\times 3-12$ configurations where both modules are connected through the same dock. After accounting for these cases, in total there are $(18\times 18\times 3-12\times 12+18\times 3-12)/2=435$ unique two module design.
@@
As the number of modules increases, interference between modules and the symmetry of the design becomes much more challenging to evaluate, so we will provide an estimation of the number of unique configurations instead. Let’s assume that there are $M$ unique configurations for an $N$ module design. When adding another module to the $N$ module design, we have $16N+2$ free docks on an $N$ module design and 18 docks on the added module, giving us a total of $(16N+2)\times 18\times 3\times M\approx 864NM$ possible configurations. Considering that swapping the added module with any module in the $N$ module design will result in an identical design, we approximately have $864NM/(N+1)$ unique designs. Using mathematical induction, we get a good estimate for the number of unique $N$ module designs to be $864^{(N-1)}/N$. This number would grow to approximately $10^{11}$ when $N=5$.
@@
To find good designs with high locomotive ability, we encoded the combinatorial configuration space into a continuous 8D latent space using a variational autoencoder (VAE; (?)). A VAE was used because it transforms the search problem into a continuous one, which simplifies optimization  (?), and because we have considerable experience using this particular encoding. We searched for good configurations within this compressed latent design space using Bayesian Optimization (BO). We used multilayer perceptrons (MLP) with rectified linear unit (ReLU) activation functions for both the VAE encoder and decoder networks. The size of each hidden layer of the VAE encoder was (512, 128, 64, 64, 16) and the size of each hidden layer of the VAE decoder was (64, 64, 128, 512).
S4.3 Representing different designs (VAE input)
To input the topologies of modules into a VAE, we represent an assembled robot as a tree, where each node is an attached module. Each configuration tree can be uniquely represented as a sequence of integers with each connection pair represented by four items: the module ID of the parent module, the position ID of the dock on the parent module, the dock ID of the child module, and the orientation ID of these two docks. In this way, an assembled design of $N$ modules can be represented by $4\times(N-1)$ integers. To encode designs with $M$ modules when $M<N$, the final $4\times(N-M)$ integers are set to a reserved value which denotes “not present”. We employed a one-hot encoding on each integer in the sequence, and then fed the concatenated result to the VAE.
The one-hot encoding represents the integer IDs as binary vectors, where only one element is set to 1 and all others are set to 0. Sixteen one-hot vectors fully specify an arbitrary configuration tree with up to five modules. Each of the one-hot vectors has 19 elements. This length ensured vectors were long enough to capture the valid range of IDs. The 16 one-hot vectors were concatenated into a single binary vector (with $16\times 19=304$ elements) and fed as input to the VAE. During reconstruction, if the decoded one-hot representation contains invalid IDs, they were replaced with the nearest valid alternative. This prevents, for example, attaching a second module to a dock already occupied by a sibling on the parent. If multiple empty tokens appear, or if an empty value occurs at an invalid position, the design was interpreted as the complete tree up to the first valid empty value.
@@
Because training time can vary greatly when evaluating configurations with differing numbers of modules, we used asynchronous Bayesian Optimization (BO), allowing multiple workers to comb the latent space for good configurations, in parallel asynchronously. We use the method in  (?) to penalize the acquisition function using information about configurations that are still under evaluation. BO was used to identify configurations that were able to locomote well. The performance of a design for the purpose of BO was taken to be the average accumulated reward across the final 10% of episodes. Candidate designs identified through this method underwent training on additional behaviors before being selected for assembly.
@@
S4.8 Advantages and limitations of the encoding
The VAE transforms the discrete space of configuration trees into a smaller, continuous latent space. This enables the use of any continuous optimization algorithm. One downside of the VAE approach, however, is that there is no guarantee that a particular design can be found or that it even exists in the latent space. For example, we were unable to identify latent coordinates corresponding to the manually designed quadruped. Passing the five-module quadrupedal configuration through the optimized encoder into latent coordinates, and then propagating those coordinates back through the optimized decoder does not fully reconstruct the original quadruped. The reconstruction contained five modules, but its forelimbs were shifted posteriorly along the spine, and its spine was elongated by an additional module, leaving a single hind limb.
Here, an important question is whether or not the compression generates latent representations that are more easily searchable than the uncompressed “raw” representation of the configuration tree. In other words, does the latent space facilitate the discovery of better designs with higher performance? To investigate this, BO was applied directly to the one-hot representation of configuration trees (see Sect. S4.3). Thus, BO directly outputs 16 probability vectors that, when converted into 16 one-hot channels, fully specify an arbitrary configuration tree with up to five modules. When using a VAE, these 16 one-hot channels were fed as input to the VAE. Without the VAE, these 16 one-hot channels were directly specified by BO. The results of this experiment are reported in Fig. S9. They suggest that BO in the latent space yields designs with higher performance than BO in the raw configuration space.
S5 Multilegged robot controllers
@@
Inspired by  (?), the reward of the walking task is a linear combination of six terms: forward velocity tracking reward $r_{\text{forward}}$, angular velocity tracking reward $r_{\text{walk-turn}}$, action rate penalty $r_{\text{action}}$, joint falling penalty $r_{\text{fall}}$, motor velocity penalty $r_{\text{motor-vel}}$, and joint acceleration penalty $r_{\text{motor-acc}}$. The forward velocity tracking reward is calculated by
@@
where $v_{\text{forward}}=\mathbf{v}\cdot\hat{\bm{x}}$ is the projection of velocity in direction $\hat{\bm{x}}$, $\mathbf{v}$ is the robot’s velocity vector, $\hat{\bm{x}}$ is the unit local forward direction vector given by the initial pose from the pose optimizer, $v^{\ast}=0.6\,\text{m/s}$ is the desired forward speed, and $\sigma_{\text{forward}}=0.15$ is a hyperparameter. The angular velocity reward is calculated by
@@
where ${\omega}^{\ast}_{z}$ is the desired angular velocity around the global z axis, which is set to 0 here, and the ${\omega}_{z}=\bm{\omega}\cdot\bm{g}_{p}$ is the z component of angular velocity in the global frame; $\bm{\omega}$ is the robot’s angular velocity and $\bm{g}_{p}$ is the projected gravity. The hyperparameter $\sigma_{\text{walk-turn}}$ is set to 0.15. The action rate penalty is calculated in the same way as training for the single module as in Eqn. S3. The joint falling penalty $r_{\text{fall}}$ is the number of joints (the sphere part of the module) touching the floor for each time step. The joint velocity penalty is calculated by
@@
which will be clipped from 0 to $1\times 10^{5}$; $\dot{{\theta}}_{i}$ is the joint velocity of joint $i$ and $\dot{\theta}_{\text{limit}}$ is the joint velocity limit whose value is 10 radians/s in our experiment. The joint acceleration penalty is calculated by
@@
Similar to the single module training, the action output by the policy, $\bm{a}_{t}$, is offset by predefined values for each joint. Before feeding the offset action to the PD controller, we filter the action with a Butterworth filter (high cutoff frequency is 3; low cutoff frequency is 0; filter order is 2) to promote the smoothness of the actions  (?). The action is then clipped to $\pm 1.2$ radians and sent to the PD controller. The torque output by the PD controller is again clipped by the TN (torque-speed) curve of the real motor.
@@
Although we have a fixed maximum episode length, we terminate each episode earlier if and when the robot falls over. Inspired by  (?), falling is determined by the deviation of the body’s upward position $d$, which is defined by:
@@
$r_{\text{pose}}=r_{\text{upward}}\,r_{\text{joint}}\\$ (S12)
where $r_{\text{upward}}$ measures how well the design stays in an upright position, defined as $r_{\text{upward}}=d$, where $d$ is the deviation of the body’s upward position defined in Eqn. S11. The pose reward $r_{\text{joint}}$ measures how close the current joint positions $\bm{\theta}$ are to the initial joint positions $\bm{\theta}_{0}$:
@@
The action rate penalty is calculated as in Eqn. S3. The weight for the pose reward term is 1 and the weight for the action rate is 0.02-0.02. The hyperparameter $\sigma_{\text{joint}}$ here is set to 10.
@@
To achieve dynamic self-righting behavior, we increase the action clipping from $\pm 1.2$ to $\pm 3.14$ radians.
@@
The observation space is the observation space for walking plus an additional dimension for a jumping command dimension, similar to  (?). The jumping command is set to 0 except when the robot is commanded to jump. In simulation, the jumping command is randomly sampled as 0 or 1 every 100 timesteps during training and is automatically reset to 0 when the robot reaches the desired height. After deploying the policy in the real world, the jump command is nominally 0 and can be set to 1 by a user. Since we assume only proprioceptive sensing on the physical hardware, rather than using a state estimator as in  (?) to measure the height of the robot for resetting the command, we set a fixed activation period $T_{\text{activate}}$ for the jumping command. $T_{\text{activate}}$ is set to 0.75 seconds in our experiment.
@@
The rewards of the jumping task consist of five terms: upward reward $r_{\text{upward}}$, pose reward $r_{\text{pose}}$, height track reward $r_{\text{height}}$, jump bonus $r_{\text{jump}}$, and turn reward $r_{\text{jump-turn}}$. The upward reward $r_{\text{upward}}$ is to prevent the robot from falling over. It is defined in the same way as in Eqn. S12. The pose reward $r_{\text{pose}}$ is to encourage the robot to retain the joint angles from the initial pose when the jumping command is not triggered. The pose reward is set to 0 when the jump signal is triggered (the jumping command in the observation is set to 1); when the jump signal is not triggered, the pose reward is defined as in Eqn. S12. The height track reward is to encourage the robot to jump to a desired height when the jumping command is triggered. When the jump signal is not triggered, $r_{\text{height}}=0$, otherwise
@@
where $h$ is the height of the torso module and $h^{\ast}$ is the desired jumping height. The values of $h^{\ast}$ for each design can be found in Table S4. To further encourage the robot to jump, a bonus $r_{\text{jump}}$ is used, which is set to 1 when the jump signal is triggered and the robot is not touching the ground; otherwise $r_{\text{jump}}=0$. The turn reward $r_{\text{jump-turn}}$ is to encourage the robot to turn when jumping. When the jump signal is triggered,
@@
where $\omega^{\ast}_{\text{jump}}=2\,\text{rad/s}$, otherwise $r_{\text{jump-turn}}$ is defined in the same way as $r_{\text{walk-turn}}$ in Eqn. S8. The reward weights are shown in Table S5.
@@
To achieve dynamic jumping actions, we increase the action clipping to $\pm 2.5$ radians from $\pm 1.2$ radians for walking.
@@
We model the amputation-agnostic control problem as a POMDP. We desire a single policy that can control a robot after arbitrary amputations. We define the state space of the robot $\mathcal{S}_{\text{meta}}$ as the collection of the state space of each module $\mathcal{S}_{\text{module}}$ in this system:
@@
where $\mathcal{S}_{\text{module}}^{i}$ is the state space of the $i$-th module and $N$ is the maximum number of modules in the system. The state space of each module includes projected gravity, angular velocity, cosine of the joint position, and the joint velocity. The action of this POMDP is the joint position of each module, and the reward function varies across the different reduced configurations (example amputations) used in training the amputation-agnostic policy (a more detailed explanation is provided below). We further reduce this POMDP to a sequence modeling problem  (?). Similar to  (?), we use a Transformer model  (?) to distill a general controller from several amputation-specific teacher controllers. However, rather than using a Transformer encoder as a student policy to directly map states to actions, we use a causally masked transformer decoder to fit a sequence of states and actions. By conditioning on the current state $\mathcal{S}_{\text{meta}}$, the transformer should predict the optimal action.
@@
For a time step $t$, we feed the most recent $K$ timesteps ($K=60$ in the experiment) of module states and actions into the Transformer, for a total of $K\times(N+1)$ tokens. Each of the module states and actions will be fed into its dedicated linear layer with layer normalization. We then add a positional embedding to the embedding of each token. Similar to Decision Transformer  (?), rather than using a standard positional embedding, we use the embedding of each timestep as the positional embedding (each of the $N+1$ tokens thus have the same positional embedding). The embeddings are then fed to a Generative Pre-trained Transformer (GPT) model  (?), which predicts the next action tokens through autoregressive modeling. The hyperparameters of the GPT model are shown in Table S9.
@@
A dataset of offline trajectories is given to train this model. In our experiment, we use the training dataset in Table S7. To generate this dataset, scenario-specific expert policies were trained from scratch to control the remnant structures. For training remnant configurations with more than one module remaining, we use the same observation space, action space, and reward function as the walking policy described above. For training a single module, we use the same observation, reward, and action as the single module rolling policy described above. However, when training against amputation examples, we add more randomization to the environment compared to the standard training methods described above. In the undamaged scenario, we randomize the length of all the limbs: Each link touching the floor is either with its nominal length or a half of its nominal length. In the four-remaining-modules scenario (one limb is removed), the farthest link of the limb diagonally across from the amputated limb is randomly assigned either its nominal length or 1.5 times its nominal length. For example, if the configuration is that the left forelimb is amputated, the right hind limb will have two possible lengths: the nominal length or an extended length. This encourages the robot to recover to an upright position. Training settings for single limb amputations of the quadruped are shown in Table S8. In the three-remaining-modules scenario (two limbs removed), we randomize the length of the forelimbs and the torso: Each link touching the ground is either with its nominal length or a half of its nominal length. When one module remains post-amputation, the left or right link is half of its nominal length. These remnant single modules were also trained on uneven terrain featuring obstacles of random heights. The heights of the obstacles are sampled from 0 to 0.03 m.
Note that the observation space of the POMDPs for generating the training dataset is different from the state space $\mathcal{S}_{\text{meta}}$. When collecting the rollouts after these policies are trained, the state of the robot $\bm{s}_{\text{meta}}\in\mathcal{S}_{\text{meta}}$ is collected rather than the observation of the specific amputation example. In addition, although the different amputation-specific policies have different zero positions for joints, we normalize them to have the same zero position when recording the rollout trajectories after training them. The training of each robot takes $1\times 10^{6}$ timesteps. To avoid the robot getting stuck at a local optimum and stopping motion, we set the weight of the forward term in the reward function as 1 at the beginning of training and tuned it down to 0.8 at the time step of $2\times 10^{5}$.